# Graph Neural Network-Based Reinforcement Learning for Controlling Biological Networks: The GATTACA Framework

### v1

@@ 6 Results
We evaluate GATTACA on the BN model environments presented in Section 5.1. The results obtained demonstrate that our approach is successful in controlling environments across a wide range of sizes.
6.1 Evaluation of iPASIP
@@
Our iPASIP method, although it is a purely statistical method based on model simulations, provides a strong alternative to traditional methods for finding BN attractor states. Table 3 presents the summary results for each model where CABEAN was able to compute the attractors for the environmental conditions specified in Table 2.
@@
To provide a more comprehensive evaluation of iPASIP, we additionally ran it on the Bladder, MAPK, T-LGL, and Bortezomib models without specifying the environmental condition – that is, simultaneously considering all of them. For comparison, we computed the attractors for each model with CABEAN in all individual environmental conditions one-by-one, i.e. in 16, 16, 128, and 32 conditions, respectively. The numbers of attractors in all environmental conditions are presented in square brackets in Table 3.
@@ F-A C-A Spurious F-A C-A Failed
Table 3: Comparison of the numbers of pseudo-attractor (PA) states identified by iPASIP with the numbers of attractors obtained with CABEAN. The numbers of fixpoint attractors (F-A), cyclic or complex attractors (C-A), and spurious pseudo-attractor states are presented for the environmental conditions in Table 2 (plain numbers) and for input nodes left unspecified, i.e. all possible environmental conditions (numbers in square brackets). The numbers of environmental conditions for which CABEAN failed to compute a result are shown in the ‘Failed’ column. The ‘?’ indicates that our verification of whether the pseudo-attractor states are indeed spurious failed.
@@
For the Bladder model, iPASIP found 25 pseudo-attractor states in total in all the 16 environmental conditions, while CABEAN returned 24 attractors in 15 environmental conditions and failed in one, namely where the EGFR_stimulus input node is set to 1 and all the remaining three input nodes are set to 0. All the 24 CABEAN-found attractors are represented within the 25 pseudo-attractor states. By computing an attractor reachable from the extra pseudo-attractor state identified by iPASIP, we found that it is in fact an attractor state of an attractor consisting of 184,320 states under the environmental condition that CABEAN was unable to handle. We further investigated the distribution of revisits to the individual attractor states. We simulated a trajectory of 1,000,000 steps with the pseudo-attractor state as the initial state. The numbers of revisits for individual states are shown in Figure 3. The extra pseudo-attractor state is clearly dominating all other attractor states in terms of the count of revisits and is correctly identified by the iPASIP method despite the attractor being unusually large.
@@
Figure 3: Bar plot of the number of revisits to individual states of the Bladder BN model attractor containing the extra pseudo-attractor state identified by iPASIP under the environmental condition of EGFR_stimulus set to 1 and all remaining input nodes set to 0. For readability, only states with at least 1,000 revisits are shown in the decreasing order of revisit count. State of index 0 is the pseudo-attractor state found by iPASIP and the remaining attractor states are indexed arbitrarily.
@@
In two cases of T-LGL and CD4+, iPASIP overestimates the number of attractor states, i.e. some of the pseudo-attractor states are spurious attractor states, although for the T-LGL model we were unable to confirm that indeed these states are spurious: we failed to compute attractors reachable from the potentially spurious states with a graph traversal approach. Nevertheless, as argued in Section 4.1, the additional states stand out due to being frequently revisited. Thus, they can still be considered viable for performing control actions. Indeed, the extra pseudo-attractor state [^1] In the spurious attractor state the nodes Cofilin, GATA3, IL4, PI3K, proliferation, AKT, PIP3_345, Dec2, IRF4, IL4RA, PDK1, and IKB are 1 and all the remaining nodes are 0. of the CD4+ BN model is spurious – this can be inferred from the fact that CABEAN determined all attractor states for the given environmental condition. By simulating 1,000,000-step trajectories from this spurious attractor state, we observed that it was revisited more than 996,000 times and the trajectories consisted of less than 70 distinct states. Clearly, in this sense, the state resembles a stable state.
@@
We verified that for all models except T-LGL all the attractors found by CABEAN have a representative amongst the pseudo-attractor states. For T-LGL, 143 of 166 CABEAN fixpoint attractors and 71 out of 130 CABEAN cyclic or complex attractors are also detected by iPASIP, some of them multiple times as more than one pseudo-attractor state may belong to the same attractor (data not shown). The remaining fixpoint pseudo-attractor states found by iPASIP are attractor states in the six environmental conditions in which CABEAN failed [^2] By failure we mean the fact that CABEAN ended with an error..
@@
In summary, the iPASIP approach provides a scalable and effective solution for identifying true pseudo-attractor states, as defined in Definition 2.1, in large-size BN models. Moreover, as the case studies show, iPASIP often successfully identifies all attractor states of a large BN model.
6.2 Evaluation of the GATTACA control framework
@@
With regard to control problem solutions found by GATTACA, the average length of the control strategies is typically within one control action/intervention of the optimal solutions computed in accordance with the formula in Equation 3 based on the control path lengths obtained by the exact ASI algorithm implemented in the CABEAN software tool, as shown in Table 4.
@@
However, for the ABA model the number of environmental conditions is too big ($2^{23}$) to be handled by CABEAN (entries marked as N/A in Table 4). As computed by GATTACA, the ABA model is characterised by a large number of pseudo-attractor states. Nevertheless, GATTACA successfully identifies effective control strategies, demonstrating the scalability of our framework in this respect.
@@
Table 4: Comparison of the average control strategy lengths required to drive the network from the source state to the target configuration using GATTACA and the exact ASI method implemented in CABEAN.
To further investigate the scalability potential of GATTACA, we analyse the MCF-7 model under all possible environmental conditions. As in the case of the ABA model, MCF-7 is characterised by a large number of pseudo-attractor states and GATTACA is capable of successfully identifying control strategies. Unfortunately, we were unable to obtain the exact results with CABEAN as it could not finish the computations of attractors within one week time on an HPC cluster node where all $256$ fully specified environmental conditions were considered.
Next, we analysed the original T-diff model with 71 nodes, introduced in [20], under two environmental conditions: 1) the original, partially specified (PS) condition from [20] and 2) the fully specified (FS) condition from [29] presented in Table 2. The former condition leaves some of the input nodes unspecified with respect to the latter, i.e. IFNGR1, IFNGR2, GP130, IL6RA, CGC, IL12RB2, IL10RB, IL10RA, IL15RA, IL2RB, and IL27RA. We refer to these conditions as the partially and fully specified environmental conditions, respectively.
The original 71-node T-diff model, under both conditions, is characterised by a large number of pseudo-attractor states identified by iPASIP, i.e. 156,864 and 25,856 states, respectively. CABEAN fails to identify the attractors and cannot compute control strategies for these BNs. However, GATTACA successfully identifies control strategies in both cases. The average control strategy length is 6.17, as shown in Table 5. It is noticeably larger than for all other BN models considered in this study. Nevertheless, this is justifiable, as such large numbers of pseudo-attractor states suggest that the two BN models have large numbers of attractors. Meanwhile, the size of the original T-diff model (71 nodes) is comparable to the size of the model in [29] (68 nodes), the latter with only 12 attractors; see Table 4. It is then highly probable that the state space of the original T-diff model under the two environmental conditions is divided into much finer-grained basins of attraction than in the case of the modified model in [29]. Consequently, driving the dynamics towards a pseudo-attractor state aligned with the given target configuration requires an action that places the network dynamics into a relatively small basin of attraction, likely necessitating simultaneous perturbation of many genes. Since this is restricted due to the limit imposed on the maximum number of genes that can be perturbed at once within a single intervention (ATGC action), identified control strategies must traverse multiple attractors to reach the target configuration.
@@
Table 5: Average control strategy lengths obtained with GATTACA (shown in the ‘GATTACA’ column) for the original 71-node T-diff model of [20] under two environmental conditions (env. cond.), i.e. the partially specified (PS) one in [20] and the fully specified (FS) one in [29]. Both conditions are characterised by large number of pseudo-attractor states identified by iPASIP.
@@
Finally, we perform one additional scalability test of GATTACA on a randomly generated BN model of 200 nodes (Random200). The model is challenging as it has no input nodes and its structure graph consists of a single SCC. Therefore, the attractor detection algorithm and the ASI control algorithm in CABEAN cannot exploit the SCC-decomposition and CABEAN fails to compute the control strategies within 168 hours (one week) on a machine with 8 Core 2.25 GHz AMD EPYC 7742 processor and 256 GB of RAM. GATTACA identifies 165 pseudo-attractor states and returns control strategies of average length of 7.1 for the target configuration $x_{3}=1$, where $x_{3}$ is one of the nodes; see Table 4. This relatively long control strategy on average can be explained by the fact that a network of highly interconnected nodes may require a simultaneous perturbation of a larger number of nodes than in the case of a loosely connected network. Similarly to the two cases of the original T-diff model in Table 5, since we limit the number of simultaneous perturbations, GATTACA finds longer control strategies that traverse a number of intermediate pseudo-attractor states.
7 Conclusions
@@
We formulated the attractor-target control problem in the context of cellular reprogramming –a novel control problem for biological networks modelled using the BN framework under the asynchronous update mode. To provide a scalable solution, we developed and implement GATTACA, a scalable computational framework based on DRL.
@@
To facilitate scalability of GATTACA, we consider our previously introduced concept of a pseudo-attractor and we further develop our PASIP procedure for effective identification of pseudo-attractor states by properly handling the situations in which model simulations get trapped in large attractors. We demonstrate the effectiveness of improved PASIP, i.e. iPASIP, on a number of BN models of real-world biological networks of different sizes by comparing the identified pseudo-attractor states with true attractors obtained with an exact state-of-the-art algorithm implemented in the CABEAN software tool. We show that iPASIP is capable of correctly identifying representative attractor states of all attractors in most of the cases. It also manages to detect attractor states in BN models for which CABEAN fails to compute the attractors. Furthermore, iPASIP proves its scalability and effectiveness as part of the GATTACA framework, where the determined pseudo-attractor states facilitate the computation of control strategies for large BN models of various real-world biological networks.
The proposed GATTACA framework offers a scalable and effective solution to the attractor-target control problem. To leverage the known structure of biological systems for encoding the network dynamics in a latent representation, we incorporate graph neural networks with graph convolutions into the artificial neural network approximator for the action-value function learned by the ATGC DRL agent. Given source attractor state and target configuration of genes, the ATGC of GATTACA finds proper control strategies that drive the network from the source pseudo-attractor to the target configuration by taking actions (intervening) only in intermediate pseudo-attractor states that correspond to phenotypical cellular states that can be observed in the lab.
@@
The results of the conducted computational experiments on a number of large real-world biological networks from literature demonstrate the scalability and effectiveness of GATTACA. Our framework is capable of finding optimal or suboptimal control strategies which lengths in most of the cases differ on average from the optimal strategies by at most one intervention. Moreover, GATTACA is capable of identifying effective control strategies of reasonable lengths in BN models on which CABEAN fails. These models are large in terms of number of nodes and often characterised by many (pseudo-)attractor states.
An additional evaluation on a challenging network which structure graph is a single large SCC of 200 nodes further demonstrates the scalability potential of GATTACA in controlling large models of complex biological networks.
8 Acknowledgments
@@
Appendix A Example
(0,0,0)(0,0,1)(0,1,0)(0,1,1)(1,0,0)(1,0,1)(1,1,0)(1,1,1)
Figure 4: Example of an STG
Let us consider a Boolean Network defined as:
$\begin{cases}x_{1}=x_{1}\\ x_{2}=x_{1}|x_{3}\\ x_{3}=x_{2}\&x_{3}\end{cases}$
It’s STG is shown in the Pic. 4. In this network $x_{1}$ is an input node, that is its state remains constant during the evolution of the network. Let us consider a target of the form $x_{2}=0$. That is we want our network to be in an attractor with $x_{2}$ set to 0, and we do not care for the state of other nodes.
If we consider the case of $x_{1}=0$ then our target is only realized by $(0,0,0)$ state, and can be controlled from the other attractor by one perturbation:
[…]
In the case of $x_{1}=1$ we have one state ralizing the target: $(1,0,0)$. It will naturally be reached via evolution from $(1,1,0)$ and we will need one perturbation to reach it from $(1,1,1)$:
[…]
Appendix B Evaluation example
We consider a BN with 5 nodes, i.e. $x_{1},x_{2},x_{3},x_{4},x_{5}$. The nodes $x_{1}$ and $x_{2}$ are designated as input nodes and the control objective is to achieve the target configuration $x_{5}=1$. Let us assume that
  • For $x_{1}=0,x_{2}=*$, the network has two attractors:
    $(0,0,0,0,0),\quad(0,0,1,1,0)$
    In both cases $x_{5}=0$, so there is no attractor aligned with the target configuration.
  • For $x_{1}=1,x_{2}=1$, the attractors are:
    $(1,1,1,0,1),\quad(1,1,0,1,0)$
  • For $x_{1}=1,x_{2}=0$, the attractor is:
    $(1,0,0,0,1)$
Thus, there are three source attractors and two target attractors.
We consider $(0,0,0,0,0)$ to be the source attractor. Without perturbations of the input nodes $x_{1}$ and $x_{2}$, the system remains trapped in the configuration $x_{5}=0$. Therefore, perturbations of input nodes are necessary to change the environmental condition. There are two options: set $x_{1}=1,x_{2}=1$ or $x_{1}=1,x_{2}=0$. We consider both.
  • Suppose that the intervention to state $(1,1,0,0,0)$ leads to attractor:
    $(1,1,0,1,0)$
    Since $x_{5}=0$, an additional intervention on $x_{3}$ is required to reach the target attractor, which is a CABEAN control path of length 1. Therefore, the total control strategy length in this case is 2.
  • The intervention to state $(1,0,0,0,0)$ leads directly to attractor:
    $(1,0,0,0,1),$
    which is a target attractor. Here, the total control strategy length is 1.
Then, the length of the reference control strategy obtained with CABEAN is 1 for this example.

### v2

@@ VII Results
We evaluate the GATTACA framework on the BN models presented in Section VI-A. The results obtained demonstrate that our approach is scalable – it is capable of controlling environments across a wide range of sizes, particularly BN model characterised by huge state spaces.
VII-A Evaluation of iPASIP
@@
Our iPASIP method, although it is a purely statistical method based on model simulations, provides a strong alternative to traditional methods for finding BN attractor states. Table III presents summarised results for each model where CABEAN was able to compute the attractors for the environmental conditions specified in Table II.
@@
To provide a more comprehensive evaluation of iPASIP, we additionally ran it on the Bladder, T-LGL, and Bortezomib models without specifying the environmental condition – that is, simultaneously considering all of them. For comparison, we computed the attractors for each model with CABEAN in all individual environmental conditions one by one, i.e. in 16, 128, and 32 conditions, respectively. The numbers of attractors in all environmental conditions are presented in square brackets in Table III.
In the case of the MAPK model, CABEAN failed in 3 of 16 possible environmental conditions. Running iPASIP under all environmental conditions resulted in a very large number of identified PA states, which we could not verify due to the lack of ground-truth information. However, running iPASIP under all environmental conditions except the ones in which CABEAN failed, resulted in fully correct identification of all attractors, see Table III.
@@ F-A C-A Spurious F-A C-A Failed
TABLE III: Comparison of the numbers of pseudo-attractor (PA) states identified by iPASIP with the numbers of attractors obtained with CABEAN. The numbers of fixpoint attractors (F-A), cyclic or complex attractors (C-A), and spurious PA states are presented for the environmental conditions in Table II (plain numbers) and for input nodes left unspecified, i.e. all possible environmental conditions (numbers in square brackets). The numbers of environmental conditions for which CABEAN failed to compute a result are shown in the ‘Failed’ column. The ‘?’ indicates that our verification of whether the PA states are indeed spurious failed.
@@
For the Bladder model, iPASIP found 25 PA states in total in all the 16 environmental conditions, while CABEAN returned 24 attractors in 15 environmental conditions and failed in one, namely where the EGFR_stimulus input node is set to 1 and all the remaining three input nodes are set to 0. All the 24 CABEAN-found attractors are represented within the 25 PA states. By computing an attractor reachable from the extra PA state identified by iPASIP, we found that it is in fact an attractor state of an attractor consisting of 184,320 states under the environmental condition that CABEAN was unable to handle.
We further investigated the distribution of revisits to the individual states of this large attractor. We simulated a trajectory of 1,000,000 steps with the PA state as the initial state. The numbers of revisits for individual states are shown in Figure 3. The extra PA state is clearly dominating all other attractor states in terms of the count of revisits and is correctly identified by the iPASIP method despite the attractor being exceptionally large.
@@
Figure 3: Bar plot of the number of revisits to individual states of the Bladder BN model attractor containing the extra PA state identified by iPASIP under the environmental condition of EGFR_stimulus set to 1 and all remaining input nodes set to 0. For readability, only states with at least 1,000 revisits are shown in the decreasing order of revisit count. State of index 0 is the PA state found by iPASIP and the remaining attractor states are indexed arbitrarily.
@@
In the cases of the T-LGL and CD4+ models, iPASIP overestimates the number of attractor states, i.e. some of the PA states are spurious attractor states, although for the T-LGL model we were unable to confirm that indeed the two states are spurious: we failed to compute attractors reachable from the potentially spurious states with a graph traversal approach. Nevertheless, as argued in Section V-A, the additional states stand out due to being frequently revisited. Thus, they can still be considered viable for performing control actions.
The extra PA state [^2] In the spurious attractor state the nodes Cofilin, GATA3, IL4, PI3K, proliferation, AKT, PIP3_345, Dec2, IRF4, IL4RA, PDK1, and IKB are 1 and all the remaining nodes are 0. of the CD4+ BN model is indeed spurious – this can be inferred from the fact that CABEAN determined all attractor states for the given environmental condition. By simulating 1,000,000-step trajectories from this spurious attractor state, we observed that it was revisited more than 996,000 times and the trajectories consisted of less than 70 distinct states. Clearly, in this sense, the state resembles a stable state.
@@
We verified that for all models except T-LGL all the attractors found by CABEAN have a representative amongst the PA states. For T-LGL, 143 of 166 CABEAN fixpoint attractors and 71 out of 130 CABEAN cyclic or complex attractors are detected by iPASIP, some of them multiple times as more than one PA state may belong to the same attractor (data not shown). The remaining fixpoint PA states found by iPASIP are attractor states in the six environmental conditions in which CABEAN failed [^3] By failure we mean the fact that CABEAN ended with an error..
@@
In summary, the iPASIP approach provides a scalable and effective solution for identifying true PA states, as defined in Definition III.1, in large-size BN models. Moreover, as the case studies show, iPASIP often successfully identifies all attractor states of a large BN model.
VII-B Evaluation of the GATTACA framework
@@
With regard to control problem solutions found by our GATTACA framework, the average length of the control strategies is typically within one control action/intervention of the optimal solution computed in accordance with the formula in Equation (2) based on the control path lengths obtained by the exact ASI algorithm implemented in the CABEAN software tool, as shown in Table IV.
@@
However, for the ABA model the number of environmental conditions is too big ($2^{23}$) to be handled by CABEAN (entries marked as N/A in Table IV). As obtained with the GATTACA framework, the ABA model is characterised by a very large number of PA states. Nevertheless, the GATTACA framework successfully identifies effective control strategies, demonstrating the scalability of our framework in this respect. Moreover, we conducted experiments in which the number of PA states was overestimated by a factor of 1000 and ATGC was still able to identify effective control strategies (data not shown). Those two facts prove that the ATGC is resilient to both high attractor counts and spurious PA states.
@@
TABLE IV: Comparison of the average control strategy lengths required to drive the network from the source state to the target configuration obtained with the GATTACA framework and the exact ASI method implemented in CABEAN. Results that could not be obtained with CABEAN are indicated as not available (N/A).
To further investigate the scalability potential of the GATTACA framework, we analysed the MCF-7 model under all possible environmental conditions. As in the case of the ABA model, MCF-7 is characterised by a large number of PA states and the GATTACA framework is still capable of successfully identifying control strategies as shown in Table V. Unfortunately, we were unable to obtain the exact results where all $256$ fully specified environmental conditions were considered with CABEAN as it could not finish the computations of attractors within 168 hours (one week) on an 8-Core 2.25 GHz AMD EPYC 7742 processor with 256 GB of RAM.
Next, we analysed the original T-diff model with 71 nodes, introduced in [36], under two environmental conditions: 1) the original, partially specified (PS) condition from [36] and 2) the fully specified (FS) condition from [25] presented in Table II. The former condition leaves some of the input nodes unspecified with respect to the latter, i.e. IFNGR1, IFNGR2, GP130, IL6RA, CGC, IL12RB2, IL10RB, IL10RA, IL15RA, IL2RB, and IL27RA.
The original 71-node T-diff model, under both conditions, is characterised by a large number of PA states identified by iPASIP, i.e. 156,864 and 25,856 states, respectively. CABEAN fails to identify the attractors and cannot compute control strategies for these BNs. However, the GATTACA framework successfully identifies control strategies in both cases. The average control strategy length is 6.17, as shown in Table V. It is noticeably larger than for all other BN models considered in this study. Nevertheless, this is justifiable, as such large numbers of PA states suggest that the two BN models have large numbers of attractors. Meanwhile, the size of the original T-diff model (71 nodes) is comparable to the size of the model in [25] (68 nodes), the latter with only 12 attractors; see Table IV. It is then highly probable that the state space of the original T-diff model under the two environmental conditions is divided into much finer-grained basins of attraction than in the case of the modified model in [25]. Consequently, driving the dynamics towards a PA state aligned with the given target configuration requires an action that places the network dynamics into a relatively small basin of attraction, likely necessitating simultaneous perturbation of many genes. Since this is restricted due to the limit imposed on the maximum number of genes that can be perturbed at once within a single ATGC action, identified control strategies must traverse multiple attractors to reach the target configuration.
@@
TABLE V: Average control strategy lengths obtained with the GATTACA framework for the MCF-7 model under all environmental conditions and the original 71-node T-diff model of [36] under two environmental conditions (env. cond.), i.e. the partially specified (PS) one in [36] and the fully specified (FS) one in [25]. Both conditions are characterised by large number of PA states identified by iPASIP.
@@
Finally, we performed one additional scalability test of the GATTACA framework on a randomly generated BN model of 200 nodes (Random200). The model is challenging as it has no input nodes and its structure graph consists of a single SCC. Therefore, the attractor detection algorithm and the ASI control algorithm in CABEAN cannot exploit the SCC-decomposition and CABEAN fails to compute the control strategies within 168 hours (one week). The GATTACA framework identifies 165 PA states and returns control strategies of average length of 7.1 for the target configuration $x_{3}=1$, where $x_{3}$ is one of the nodes; see Table IV. This relatively long control strategy on average can be explained by the fact that a network of highly interconnected nodes may require a simultaneous perturbation of a larger number of nodes than in the case of a loosely connected network. Similarly to the two cases of the original T-diff model in Table V, since we limit the number of simultaneous perturbations, the GATTACA framework finds longer control strategies that traverse a number of intermediate PA states.
VIII Conclusions
@@
We formulate the attractor-target control problem in the context of cellular reprogramming – a novel control problem for biological networks modelled using the BN framework under the asynchronous update mode. To provide a scalable solution, we develop and implement the GATTACA framework. Our framework can handle models of sizes that are prohibitive for the existing state-of-the-art control algorithms.
To facilitate scalability, the GATTACA framework is based on DRL. Importantly, to leverage the known structure of a biological system and to encode the network dynamics into a latent representation, we incorporate graph neural networks with graph convolution operations into the artificial neural network approximator for the action-value function learned by the ATGC DRL agent. The available knowledge of the structure of a BN model is explicitly incorporated into the ATGC by defining the GCN architecture in accordance with the structure graph of the BN. Through interactions with the BN behaviour determined by the Boolean functions of the BN, the BN model dynamics is implicitly encoded into the ATGC via training. Moreover, the GCN creates an embedding of a discrete graph into a continuous space, which enhances the training of the DRL agent.
@@
Moreover, the scalability is achieved by considering previously introduced concept of a pseudo-attractor and by exploiting iPASIP, which we develop by improving previously proposed approach for effective identification of PA states in large BN models. We demonstrate the effectiveness of iPASIP on a number of BN models of real-world biological networks of different sizes by comparing the identified PA states with true attractors obtained with an exact state-of-the-art algorithm implemented in the CABEAN software tool. We show that iPASIP is capable of correctly identifying representative attractor states of all attractors in most of the cases. It also manages to detect attractor states in BN models for which CABEAN fails to compute the attractors. Furthermore, iPASIP proves its scalability and effectiveness as part of the GATTACA framework, where the determined PA states facilitate the computation of control strategies for large BN models of various real-world biological networks.
@@
The proposed GATTACA framework offers a scalable and effective solution to the attractor-target control problem as demonstrated by he results of the conducted computational experiments on a number of large real-world biological networks from literature. Given source attractor state and target configuration of genes, the ATGC of the GATTACA framework finds proper control strategies that drive the network from the source pseudo-attractor to the target configuration by taking actions (intervening) only in intermediate PA states that correspond to phenotypical cellular states that can be observed in the lab. Our framework is capable of finding optimal or suboptimal control strategies which lengths in most of the cases differ on average from the optimal strategies by at most one intervention. Moreover, the GATTACA framework is capable of identifying effective control strategies of reasonable lengths in BN models on which CABEAN fails. These models are large in terms of number of nodes and often characterised by many (pseudo-)attractor states. An additional evaluation conducted on a challenging network which structure graph is a single large SCC of 200 nodes, which analysis is beyond the capabilities of state-of-the-art exact solutions, further demonstrates the unprecedented scalability potential of the GATTACA framework in controlling large models of complex biological networks.
IX Acknowledgments